OpenStudy (anonymous):

give exactvalue if possible for sin(60+x)-cos(30+x)

OpenStudy (campbell_st):

several ways to do this here is 1. well you need the sum expansions \[\sin(60 + x) - \cos(30 + x) = \sin(60)\cos(x) + \cos(60)\sin(x) - \cos(30)(\cos(x) + \sin(30)\sin(x)\] substitute the exact values for sin(60), cos(60) sin(30) and cos(30) then simplify hope this helps

OpenStudy (anonymous):

i did and does the answer become zero?

OpenStudy (campbell_st):

I got sin(x)